OpenStudy (anonymous):

What is the vertex of the following inequality: 2(2x^2 – 3x) > -9 ?i have to pick one of them for the answer (-3/4, 27/4) (3/4, -27/4) (-3/4, -27/4) (3/4, 27/4)

OpenStudy (anonymous):

help

OpenStudy (nurali):

2(2x^2 – 3x) > -9 4x^2-6x>-9 4x^2-6x+9>0 f(x) = ax^2 + bx + c The vertex = (-b/2a, f(-b/2a)) Find the vertex of f(x) = 4x^2 - 6x + 9 -b/2a=6/2*4 -b/2a=6/8 -b/2a=3/4 f(x) = 4x^2 - 6x + 9 f(3/4) = 4(3/4)^2 - 6(3/4) + 9 f(3/4)=4(9/16)-18/4+9 f(3/4)=9/4-18/4+9 f(3/4)=9-18+36/4 f(3/4)=27/4 So, the vertex is (3/4,27/4).

OpenStudy (nurali):

Welcome to Openstudy @juanaandrea

OpenStudy (anonymous):

wow thank you.... you just explained everything ur great

OpenStudy (nurali):

My Pleasure.

OpenStudy (anonymous):

ok i will thanks

OpenStudy (nurali):

Anytime.